Calculate the molar solubility of Al(OH)in a solution that is slightly basic with a pH of 9.50.

Enter your response in scientific notation, e.g. enter 2E3 for 2000.


1
Expert's answer
2022-04-11T22:28:04-0400

Al(OH)3 = Al3+ + 3OH-

Ksp[Al(OH)3] = [Al3+] [OH-]3 = 5.0 × 10-33

[OH-] = 10-pOH = 10-(14-pH) = 10-(14-9.5) = 10-4.5 = 3.16 × 10-5 M

[Al3+] [OH-]3 = S × (3.16 × 10-5)3 = 5.0 × 10-33

S = 1.58 × 10-19 M